Three charges lie in the plane in a vacuum. The distance between q1 and q2 is 5.0 cm while the distance from q2 and q3 is 7.5cm. If q1 = +3.0 nC, q2= -1.5 nC and q3= +2.0 nC, find the magnitude and direction of the net electrostatic force on q2. Note that 1nC = 1x10-9C

1
Expert's answer
2021-02-08T18:41:02-0500

Since the second charge is between the other two, it is subjected to the action of two forces. The first charge attracts it, the third charge repel it, which means that the force on the second charge is toward charge 1. Calculate the magnitude:


"F=F_1+F_2,\\\space\\
F=k\frac{q_1q_2}{r_{12}^2}+k\frac{q_3q_2}{r_{23}^2}=\\\space\\
F=kq_2\bigg(\frac{q_1}{r_{12}^2}+\frac{q_3}{r_{23}^2}\bigg)=\\\space\\
=-2.1\cdot10^{-5}\text{ N}."